OpenStudy (the.lit.owl_23):

would somebody please give me the answer of help me out :) "I think the answer is B" How many solutions does the following equation have? |4x + 12| = 0 A. No solution B. One solution C. Two solutions D. Infinitely many solutions

OpenStudy (avi13):

correct

OpenStudy (the.lit.owl_23):

u think so?

OpenStudy (the.lit.owl_23):

hello?

OpenStudy (avi13):

no it is x=-3

OpenStudy (the.lit.owl_23):

so it still is "one solution"

OpenStudy (the.lit.owl_23):

umm?

OpenStudy (the.lit.owl_23):

..I guess I am going with B?

OpenStudy (avi13):

ohhh sorry i mean yes one soln

OpenStudy (avi13):

B

OpenStudy (the.lit.owl_23):

ok thx

OpenStudy (slickster256):

4 * -3 + 12 = 0 -12 + 12 = 0 0 = 0

OpenStudy (the.lit.owl_23):

no solution ?

OpenStudy (the.lit.owl_23):

@slickster256 ???????????///

OpenStudy (the.lit.owl_23):

welp thanks for the help :)

OpenStudy (the.lit.owl_23):

I am gonna close the question now ..

OpenStudy (slickster256):

sorry, there is only one (i plugged -3 which was avi's answer into it to show u it was correct